OpenStudy (anonymous):

Which equation represents the graph of the linear function? A. y = 2 B. y = −2 C. y = 2x D. y = −2x

OpenStudy (austindeboe18):

The answer is C

OpenStudy (anonymous):

did you even look at the graph

OpenStudy (anonymous):

@jakashaka123

OpenStudy (austindeboe18):

I didn't know there was a graph hang on

OpenStudy (austindeboe18):

im sorry its A. y=2

OpenStudy (anonymous):

its ok thx man

OpenStudy (austindeboe18):

welcome

OpenStudy (anonymous):

Complete the equation of the graphed linear function. Write the slope in decimal form. y = x + The graph of the function goes through point 0, negative 1 and point 2, 0.

OpenStudy (anonymous):

@austindeboe18

OpenStudy (austindeboe18):

y = 4/5x + 4, And 4/5 = 0.8. y = 0.8x + 4.

OpenStudy (anonymous):

y=1/2x-1

OpenStudy (anonymous):

so 0.8x +4

OpenStudy (austindeboe18):

yes

OpenStudy (anonymous):

Slope intercept is y=mx+b m=slope b=intercept and yes ...

OpenStudy (anonymous):

omplete the equation of the graphed linear function in point-slope form. y – (–2) = (x – ) The graph of the function goes through point 1, negative 2 and point 2, 2.

OpenStudy (anonymous):

@austindeboe18

OpenStudy (anonymous):

@i_luv_donuts

OpenStudy (austindeboe18):

one sec @R3DN3CKR3B3L187

OpenStudy (anonymous):

ok @austindeboe18

OpenStudy (austindeboe18):

y= 2x+18